OpenStudy (superman499):

Solve for x. -3 + 8x - 5 = -8

DivineSolar (divinesolar):

Hold on please :)

DivineSolar (divinesolar):

Let's solve your equation step-by-step. −3+8x−5=−8 Step 1: Simplify both sides of the equation. 8x−8=−8 Step 2: Add 8 to both sides. 8x−8+8=−8+8 8x=0 Step 3: Divide both sides by 8. 8x 8 = 0 8 x=0 Any questions?

DivineSolar (divinesolar):

Refresh your page.

DivineSolar (divinesolar):

Any questions or is this good enough ^-^!

DivineSolar (divinesolar):

@Superman499 You there?